# Imaging Shapes of Atomic Nuclei in High-Energy Nuclear Collisions

### v1

@@ Methods
The total systematic uncertainties are combined in quadrature from different sources, and in some cases, they are larger than the statistical uncertainties. The total uncertainties of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are 2.5–4%, 2–5%, and 4–10%, respectively. Note that for the ratios between U+U and Au+Au, their uncertainties are evaluated for each source and combined in quadrature to form the total systematic uncertainties. This process results in a partial cancellation of the uncertainties between the two systems.
.9 Assigning uncertainties on $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$
@@
Here, $\bar{\beta}_{a}=0.294$ and $\sigma_{a}=0.021$ represent the mean and uncertainty of $\beta_{2\mathrm{U}}$ extracted from $R_{(\delta p_{\mathrm{T}})^{2}}$ in Fig. 3g. Similarly, $\bar{\beta}_{b}$ and $\sigma_{b}$ are the mean and uncertainty of $\beta_{2\mathrm{U}}$ from $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, and they depend on the parameter $\gamma_{\mathrm{U}}$. We sample a uniform prior distribution in $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$ to obtain the posterior distribution. From this posterior distribution, we obtained the mean and one standard deviation uncertainty of $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$, as well as the confidence contours displayed in Fig. 3g.
We also performed an extraction using all three ratio observables. In this case, an additional Gaussian probability density function derived for $R_{v_{2}^{2}}$ with a mean $\bar{\beta}_{c}=0.240$ and variance $\sigma_{c}=0.014$ is multiplied to Eq. (10), and the same procedure described above is repeated. This gives $\beta_{2\mathrm{U}}=0.281\pm 0.010$ and $\gamma_{\mathrm{U}}=4.4^{\circ}\pm 3.2^{\circ}$. However, we do not use this estimate, given the imperfect description of $v_{2}$ in the hydrodynamic model.
.10 Hydrodynamic model setup and simulation
Table 1 details the Woods-Saxon parameters for Au and U used in the IP-Glasma+MUSIC model calculations. The nucleon-nucleon inelastic cross-sections are chosen to be the standard values 42 mb and 40.6 mb for Au+Au collisions at 200 GeV and U+U collisions at 193 GeV, respectively. For U, the nuclear shape in Eq. (1) is extended to include a possible small axial hexadecapole deformation $\beta_{4}$:
Most low-energy nuclear structure models favor a modest oblate deformation for ${}^{197}$u Bally et al. (2023) in the range of $\beta_{2\mathrm{Au}}\approx 0.12-0.14$. We assume ${}^{197}$u is oblate shape with $\beta_{2\mathrm{Au}}=0.14$ as the default choice for simulation. For ${}^{238}$, we have scanned several $\beta_{2\mathrm{U}}$ values ranging from 0 to 0.34. We also vary $\beta_{4\mathrm{U}}$ from 0 to 0.09 and $\gamma_{\mathrm{U}}$ in the range of ${}^{\circ}$${}^{\circ}$ to examine the sensitivity of the U+U results to hexadecapole deformation and triaxiality. For each setting, about 100,000–400,000 events are generated using the officially available IP-Glasma+MUSIC Schenke et al. (2020a, b). Each event is oversampled at least 100 times to minimize non-flow effects and statistical fluctuations in the hadronic transport. These calculations were performed using services provided by the Open Science Grid (OSG) Consortium Pordes et al. (2007); Sfiligoi et al. (2009).
The role of final state effects is studied by varying the shear and bulk viscosities simultaneously up and down by 50%. The impact on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are shown for Au+Au collisions in top of panels of Fig. 6. The values of these flow observables are changed by more than a factor of two as a function of centrality, yet the ratios between U+U and Au+Au collisions, shown in the bottom panels, are relatively stable. A small reduction of $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ are observed in the non-central collisions, when values of viscosities are halved. However, this change is clearly an overestimate since the calculated flow observables greatly overestimate the data. So in the end, half of the variations of the ratios are included in the model uncertainty.
The main theoretical uncertainties arise from variations in nuclear structure parameters. Parameters common between two collision systems, such as the minimum inter-nucleon distance in nuclei $d_{\mathrm{min}}$, are not expected to contribute to the uncertainty significantly. However, other parameters including nuclear radius $R_{0}$, skin $a$, and higher-order deformation $\beta_{4}$ could be different between Au and U, and hence contribute more to the theoretical uncertainty.
Table 1 provides a list of variations of nuclear structure parameters. The impact of these variations on ratios of flow observables is displayed in Fig. 7. The ratios of flow observables are insensitive to these variations in the most central collisions. Clearly, $R_{v_{2}^{2}}$ is particularly sensitive to skin parameter $a$. This is understandable, as $v_{2}$ has a large contribution from the reaction plane flow, which varies strongly with the value of $a$ Jia et al. (2023).
Model uncertainties for the ratios are derived by combining the impact of varying viscosities, together with various sources from Fig. 7 according to the Barlow prescription Barlow (2002). As a consequence, checks that are consistent with the default calculation within their statistical uncertainties do not contribute to the model uncertainties. The combined model uncertainties for one standard deviation are illustrated in Fig. 3.
A cross-check was also carried out for an alternative hydrodynamic code, the Trajectum model Nijs and van der Schee (2023); Giacalone et al. (2023). This model has 20 parameter sets obtained from a Bayesian analysis of the Pb+Pb data at the LHC, but was unfortunately not tuned to the RHIC data. For this calculation, we simply repeat the calculation at RHIC energy and calculate the same observables. While the description of the $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ is reasonable, several parameter sets give negative values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in mid-central collisions, and subsequently not used. The calculation is performed for the remaining 16 parameter sets as a function of centrality, and root mean square variation among these calculations is assigned as the uncertainty.
Figure 8 displays the ratios of flow observables from Trajectum and compares them to IP-Glasma+MUSIC. The results from these two models agree within their uncertainties for $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$, with Trajectum predictions slightly higher in the UCC region. This leads to slightly lower values of $\beta_{2\mathrm{U}}$: $\beta_{2\mathrm{U}}=0.230\pm 0.010$ for $R_{v_{2}^{2}}$ and $\beta_{2\mathrm{U}}=0.273\pm 0.015$ for $R_{(\delta p_{\mathrm{T}})^{2}}$.
For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, however, the Trajectum model tends to under-predict the data. In central collisions, such discrepancy can be improved by adjusting the triaxiality parameter to $\gamma_{\mathrm{U}}=15^{\circ}$. Overall, the comparison of the Trajectum model with data leads to similar constraints on $\beta_{2\mathrm{U}}$ but slightly larger $\gamma_{\mathrm{U}}$.
Table 1: The choices of Woods-Saxon parameters in Eqs. (3) and (11), including deformations, in the IP-Glasma+MUSIC model. The default values are denoted by bold font, while the rest are variations designed to constrain the values of $(\beta_{2\mathrm{U}},\gamma_{\mathrm{U}})$ and derive theoretical uncertainties associated with other structure parameters.
Species R(fm) a (fm) $d_{\mathrm{min}}$(fm) $\beta_{2}$ $\beta_{4}$ $\gamma$ (${}^{\circ}$)
${}^{197}$u 6.62,6.38 0.52 0.9, 0.4 0.14,0.12 0 60
${}^{238}$ 6.81,7.07 0.55 0.9, 0.4 0,0.15,0.22 0.09,0 0, 10, 15
0.495,0.605 0.25,0.28,0.34 20

### v2

@@ Methods
The default results are obtained from the two-subevent method. The total systematic uncertainties, including these sources and non-flow, are calculated as a function of centrality. The uncertainties of the ratios between U+U and Au+Au are evaluated for each source and combined in quadrature to form the total systematic uncertainties. This process results in a partial cancellation of the uncertainties between the two systems. The uncertainties from different sources discussed above on the ratios are shown by the black boxes in Fig. 6. The total systematic uncertainties including non-flow in the 0–5% centrality range amount to 3.9%, 4.4%, and 12.5% for $R_{v_{2}^{2}}$, $R_{(\delta p_{\mathrm{T}})^{2}}$, and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
.10 Hydrodynamic model setup and simulation
Table 1 details the Woods-Saxon parameters for Au and U used in the IP-Glasma+MUSIC model calculations. The nucleon-nucleon inelastic cross-sections are the standard values 42 mb and 40.6 mb for Au+Au collisions at 200 GeV and U+U collisions at 193 GeV, respectively. For U, the nuclear shape in Eq. (1) is extended to include a possible small axial hexadecapole deformation $\beta_{4}$:
Most low-energy nuclear structure models favor a modest oblate deformation for 197Au Bally et al. (2023). We assume $\beta_{2\mathrm{Au}}=0.14$ and and $\gamma_{\mathrm{Au}}=45^{\circ}$ as the default choice for 197Au, which are varied within the range in the range of $\beta_{2\mathrm{Au}}\approx 0.12-0.14$ and $\gamma_{\mathrm{Au}}\approx 37-53^{\circ}$ according to Refs. Ryssens et al. (2023); Bally et al. (2023). These calculations reasonably reproduce many observables related to the ground-state nuclear deformation. For 238U, we scan several $\beta_{2\mathrm{U}}$ values ranging from 0 to 0.34. We also vary $\beta_{4\mathrm{U}}$ from 0 to 0.09 and $\gamma_{\mathrm{U}}$ in the range of 0–20 to examine the sensitivity of the U+U results to hexadecapole deformation and triaxiality. For each setting, about 100,000–400,000 events are generated using the officially available IP-Glasma+MUSIC Schenke et al. (2020a, b). Each event is oversampled at least 100 times to minimize statistical fluctuations in the hadronic transport. These calculations were performed using services provided by the Open Science Grid (OSG) Consortium Pordes et al. (2007); Sfiligoi et al. (2009).
The role of final state effects is studied by varying the shear and bulk viscosities simultaneously up and down by 50%. The impacts on $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are shown for Au+Au collisions in top of panels of Fig. 7. The values of these flow observables are changed by more than a factor of two as a function of centrality. Yet, the ratios between U+U and Au+Au collisions, shown in the bottom panels, are relatively stable. A small reduction of $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ are observed in non-central collisions, when values of viscosities are halved. However, this change is an overestimate since the calculated flow observables greatly overestimate the data. So, in the end, half of the variations of the ratios are included in the model uncertainty.
Fig. 7: Sensitivity to shear and bulk viscosities of the QGP. IP-Glasma+MUSIC model prediction of the $\left\langle v_{2}^{2}\right\rangle$ (left), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (middle) and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (right) compared with the data in Au+Au collisions and the corresponding ratios (bottom) between U+U and Au+Au collisions, for different amount of shear and bulk viscosities.
The main theoretical uncertainties arise from variations in nuclear structure parameters. Parameters common between two collision systems, such as the minimum inter-nucleon distance in nuclei $d_{\mathrm{min}}$, are not expected to contribute to the uncertainty significantly. However, other parameters, including nuclear radius $R_{0}$, skin $a$, and higher-order hexadecapole deformation $\beta_{4}$, could be different between Au and U and hence contribute more to the theoretical uncertainty.
Table 1 provides a list of variations of nuclear structure parameters. The impact of these variations on ratios of flow observables is displayed in Fig. 8. The ratios of flow observables are insensitive to these variations in the most central collisions. $R_{v_{2}^{2}}$ is particularly sensitive to skin parameter $a$. This is understandable, as $v_{2}$ has a large contribution from the reaction plane flow, which varies strongly with the value of $a$ Jia et al. (2023).
Fig. 8: Sensitivity to nuclear structure parameters. IP-Glasma+MUSIC model prediction of the ratios of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) between U+U and Au+Au collisions. The calculations are done for different Glauber model parameters, and they are compared with the data.
Fig. 9: Comparison between two hydrodynamic models. The ratios of $\left\langle v_{2}^{2}\right\rangle$ (left), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (middle) and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (right) as a function of centrality from IP-Glasma+MUSIC (solid lines) and Trajectum (dashed lines), assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue). For the IP-Glasma+MUSIC model, only the uncertainties for the default configuration are shown for clarity. They are compared with the data.
Model uncertainties for the ratios are derived by combining the impact of varying viscosities, together with various sources from Fig. 8. As a consequence, checks that are consistent with the default calculation within their statistical uncertainties do not contribute to the model uncertainties. The combined model uncertainties for one standard deviation are displayed in Fig. 3.
A cross-check is conducted for an alternative hydrodynamic code, the Trajectum model Nijs and van der Schee (2023); Giacalone et al. (2023). This model has 20 parameter sets obtained from a Bayesian analysis of the Pb+Pb data at the LHC, but was not tuned to the RHIC data. For this calculation, we simply repeat the calculation at RHIC energy and calculate the same observables. While the description of the $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ is reasonable, several parameter sets give negative values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in mid-central collisions, and subsequently not used. The calculation is performed for the remaining 16 parameter sets as a function of centrality, and root mean square variations among these calculations are assigned as the uncertainty.
Figure 9 displays the ratios of flow observables from Trajectum and compares them to IP-Glasma+MUSIC. The results from these two models agree within their uncertainties for $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$, with Trajectum predictions slightly higher in the UCC region. This leads to slightly lower values of $\beta_{2\mathrm{U}}$ than the IP-Glasma model: $\beta_{2\mathrm{U}}=0.228\pm 0.013$ for $R_{v_{2}^{2}}$ and $\beta_{2\mathrm{U}}=0.276\pm 0.018$ for $R_{(\delta p_{\mathrm{T}})^{2}}$.
For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, however, the Trajectum model tends to systematically under-predict the data, as well as has much larger uncertainties compared with IP-Glasma model. In central collisions, such discrepancy can be improved by using a larger triaxiality parameter value $\gamma_{\mathrm{U}}\sim 15^{\circ}$. Overall, the comparison of the Trajectum model with data gives similar constraints on $\beta_{2\mathrm{U}}$ with comparable uncertainties but larger $\gamma_{\mathrm{U}}$ value with bigger uncertainties (next section).
Table 1: Choices of Woods-Saxon parameters in Eqs. (3) and (11), including deformations, in the IP-Glasma+MUSIC model. The default values are denoted by bold font, and the rest are variations designed to constrain the values of $(\beta_{2\mathrm{U}},\gamma_{\mathrm{U}})$ and derive theoretical uncertainties associated with other structure parameters.
Species R(fm) a (fm) $d_{\mathrm{min}}$(fm) $\beta_{2}$ $\beta_{4}$ $\gamma$ ()
197Au 6.62,6.38 0.52 0.9, 0.4 0.14,0.12 0 53, 45, 37
238U 6.81,7.07 0.55 0.9, 0.4 0,0.15 0.09,0 0, 10
0.495 0.22, 0.25 15, 20
0.605 0.28,0.34
.11 Assigning uncertainties on $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$
@@
Here, $\bar{\beta}_{a}=0.294$ and $\sigma_{a}=0.021$ represent the mean and uncertainty of $\beta_{2\mathrm{U}}$ extracted from $R_{(\delta p_{\mathrm{T}})^{2}}$ in Fig. 3g from the IP-Glasma+MUSIC model. Similarly, $\bar{\beta}_{b}$ and $\sigma_{b}$ are the mean and uncertainty of $\beta_{2\mathrm{U}}$ from $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, and they depend on the parameter $\gamma_{\mathrm{U}}$. We sample a uniform prior distribution in $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$ to obtain the posterior distribution. From this posterior distribution, we obtained the mean and one standard deviation uncertainty of $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$, $\beta_{2\mathrm{U}}=0.297\pm 0.015$ and $\gamma_{\mathrm{U}}=8.5^{\circ}\pm 4.8^{\circ}$, as well as the confidence contours displayed in Fig. 3g. This statistical analysis is also performed for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ for the Trajectum model, yielding constraints of $\beta_{2\mathrm{U}}=0.275\pm 0.017$ and $\gamma_{\mathrm{U}}=15.5^{\circ}\pm 7.8^{\circ}$.
Finally, we perform an analysis to combine the constraints of the IP-Glasma+MUSIC and Trajectum models. This is achieved by multiplying the probability density function Eq. (12) from the two models, treating their constraints as statistically independent. This approach yields $\beta_{2\mathrm{U}}=0.286\pm 0.012$ and $\gamma_{\mathrm{U}}=8.7^{\circ}\pm 4.5^{\circ}$. We noticed that Trajectum model does not impact the constraints on $\gamma_{\mathrm{U}}$ due to the model’s large uncertainty, but the uncertainty on $\beta_{2\mathrm{U}}$ reduces significantly due to the comparable precision in the two models. Therefore, we also include the difference of the extracted $\beta_{2\mathrm{U}}$ values between the two models as an additional theoretical uncertainty. The final constraints given by this procedure are $\beta_{2\mathrm{U}}=0.286\pm 0.025$ and $\gamma_{\mathrm{U}}=8.7^{\circ}\pm 4.5^{\circ}$.
.12 Data availability
All raw data for this study were collected using the STAR detector at Brookhaven National Laboratory and are not available to the public. Derived data supporting the findings of this study are publicly available in the HEPData repository (https://www.hepdata.net/record/147196) or from the corresponding author on request.
.13 Code availability
The codes to process raw data collected by the STAR detector are publicly available on GitHub (https://github.com/star-bnl). The codes to analyse the produced data are not publicly available.